OpenStudy (anonymous):

Explain, in complete sentences, the difference between a radical equation and an equation simply containing a radical expression. Provide an example of each.

OpenStudy (anonymous):

no

OpenStudy (dmezzullo):

@ryan123345

OpenStudy (anonymous):

Can you tell me what does a radical mean?

OpenStudy (anonymous):

its like the \[\sqrt[3]{6}^2\]

OpenStudy (anonymous):

Right, so in an equation where the variable is a radical, like x^(1/2) +1 =0 is a radical equation. And an equation that has a radical is a simple equation, like x+ 2^(1/2)=0

OpenStudy (anonymous):

An expression does not contain an equals sign. This holds true for all expressions, not just ones with radicals. They look like this. (sqrt(x)+17). If you get one of these as a question, they've most likely given you x. An equation has an equals sign in it. With these, you're almost always supposed to find the value of x. They look like this. (sqrt(x)=16). Hope that's what you were looking for!

OpenStudy (anonymous):

@singlesixx I love you. That is all

OpenStudy (anonymous):

Haha you're very welcome.